OpenStudy (anonymous):

Two complementary angles measure (2x+10)° and (x+10)° what the value of x ?

OpenStudy (aripotta):

complementary angles means that, when added together, they equal 90 degrees. so (2x + 10) + (x + 10) = 90 solve for x

OpenStudy (anonymous):

X= 20?????

OpenStudy (aripotta):

2x + 10 + x + 10 = 90 3x + 20 = 90 3x = 70 x = 70/3 or 23 1/3

OpenStudy (anonymous):

X=23

OpenStudy (anonymous):

?

OpenStudy (aripotta):

23 and 1/3

OpenStudy (anonymous):

Thanks